OpenStudy (anonymous):

solve the following system 4x-y=-3 x=2y-13

OpenStudy (anonymous):

x=1, y=7 medal please!!!!!!!!!!!!!!!!!!!!!

OpenStudy (katrinakaif):

4(2y - 13) - y = -3 8y - 52 - y = -3 7y = 49 y = 7 Now plug into the orginal equation to get x =)

OpenStudy (anonymous):

can u help me with this one... solve the system 8x+2y=-3 y=-4x